OpenStudy (anonymous):

what is the function of cos that has an amplitude of 4 period of 2x and has a point at (0,2)

OpenStudy (anonymous):

the amplitude should be 4, therefore it should be 4cosx. period should be 2x... this doesn't quite make sense I think the question is that the period should be 2pi

OpenStudy (anonymous):

So how do i answer the question?

OpenStudy (ybarrap):

If you want an amplitude of A and a period of \(T\), then the equation you want is $$ A\cos(\cfrac{2\pi}{T}t) $$ Where \(t\) is the independent variable. Now, to make this function go through <0,2>, use the phase, \(\theta\): $$ y=A\cos(\cfrac{2\pi}{T}t+\theta) $$ The objective then is to find the phase such that y passes through <0,2>: $$ 2=A\cos(\cfrac{2\pi}{T}(0)+\theta)\\ 2=A\cos(\theta)\\ \implies \theta = \arccos \cfrac{2}{A} $$ That's it! Make sense?